# A Survey on Adversarial Contention Resolution

### v1

@@ 4. Jamming-Resistant Contention Resolution
There are a number of papers addressing applied security considerations with respect to jamming (for examples, see (bayraktaroglu:performance, ; xu:feasibility, ; lin:link, ; aschenbruck:simulative, ; wood:denial, ; xu:jamming, ; deejam:wood, ), and surveys (mpitziopoulos:survey, ; VADLAMANI201676, )). For example, spread spectrum techniques involve devices coordinating their communication over different portions of the spectrum in order to evade collisions (spread:foiling, ; navda:using, ). Another defense is to have devices map the areas impacted by jamming and then routing traffic around those areas (wood2003jam, ; han2018hierarchical, ). Specialized antenna technology (i.e., controlled reception pattern antennas) can be used to filter out a jamming signal, which is a process referred to as null steering (zheng2018design, ; mcmilin2015gps, ).
In terms of communication on a multiple access channel, there is a large body of results showing that several fundamental distributed computing challenges—such as broadcast (koo2, ; gilbert:near, ; gilbert:making, ; chen:broadcasting, ), leader election (gilbert:malicious, ; gilbert2009malicious, ), gossip (gilbert:interference, ; dolev:gossiping, ), node discovery (meier:speed, ), and simple point-to-point communication (king:conflict, ; DBLP:journals/dc/KingPSY18, )—can be solved despite a powerful jamming adversary; and many of these results have been surveyed previously (young:overcoming, ; Bender:2015:RA:2818936.2818949, ). Here, we focus on prior contention-resolution results that offer provable guarantees on throughput or makespan, despite a powerful jamming adversary.
@@
A jamming adversary poses at least two significant new challenges over the dynamic-arrivals model considered in Section 3. First, it may not be obvious that non-trivial makespan guarantees are possible; for example, if jamming occurs in every slot in perpetuity, then no participants can succeed. Therefore, we must either constrain the jamming power of the adversary (for example, by assuming a bounded adversary), or alter our measure of throughput to address this case. Second, the algorithms surveyed so far that rely on channel feedback to adjust contention—for example, by estimating $n$ (such as in (BenderKPY18, ; DBLP:journals/jacm/GreenbergFL87, )) or by exchanging messages between participants—are vulnerable, since jamming allows the adversary to create collisions and prevent communication.
With regard to the first challenge, Chang et al. (ChangJP19, ) measure performance over non-jammed slots. Specifically, assuming that at least one participant is present in the system, a slot that is not jammed, but where no participant succeeds, is said to be wasted. Guaranteeing that only a small constant fraction of slots are wasted allows us to derive meaningful results despite an adversary that jams aggressively, perhaps even in every slot.
Tuning the contention requires care. Participants follow some simple rules for updating the sending probability in each slot: increasing if there is silence, decreasing if there is noise, and leaving it unchanged in the case of a success. The adversary may lower the sending probabilities of participants by jamming (i.e., adding noise), thus depressing contention. However, the update rule allows for contention to quickly recover once the jamming ceases (although, it may resume again later), and few slots are wasted as a result. Notably, this design allows the algorithm to achieve roughly $1/e$ throughput in the non-jammed slots, which rivals that of slotted ALOHA, despite the more-challenging adversarial setting.
@@
An early series of results against a bounded adversary was obtained by Awerbuch et al. (awerbuch:jamming, ), who consider the following model. There are $n$ honest participants equipped with CD, and the network is saturated (recall Section 2.1). For some $T\in\mathbb{N}$ and constant $10<\epsilon<$ 1, there is an adversary that can jam at most $(1-\epsilon)w$ of the slots in any consecutive $w\geq T$ slots. This adversary is referred to as $(T,1-\epsilon)$-bounded, and the value of $\epsilon$ is unknown to participants.
@@
Such bounded jamming may be well-motivated in practice, where aggressive jamming can reveal that an attack is occurring, as well as the location of the malicious device(s) (for examples, see (liu2009localizing, ; liu2013error, ; xu:feasibility, )), or correspond to a limited power supply. The adversary is also adaptive, meaning that the adversary knows the full history of the actions performed by honest nodes for all the previous slots, and may use this information to inform its jamming strategy in the current slot; however, the adversary does not know what actions participants are taking in the current slot. When the adversary jams a slot, all listening participants hear that the slot is jammed; this is referred to as a 1-uniform adversary.
@@
Awerbuch et al. (awerbuch:jamming, ; DBLP:journals/talg/AwerbuchRSSZ14, ) define an algorithm to be constant competitive if, for any sufficiently large number of slots, the honest nodes perform successful message transmissions in a constant fraction of the non-jammed steps. The authors provide a randomized algorithm that is constant competitive with high probability (w.h.p.) in $n$, when all participants know a common parameter, $\gamma$$=O(1/(\log T+\log\log n))$, and the algorithm is run for a sufficiently large number of slots (as a function of: $T$, polylog($\max\{T,n\}$)), and $1/\gamma$). Roughly speaking, in a slot, the algorithm dictates that each participant $u$ sends a message with probability $p_{u}$. If $u$ does not send, it listens and updates $p_{u}$ via a multiplicative factor which can increase or decrease, depending on whether or not the slot is idle. Notably, this multiplicative weights update method shares common ground with the later algorithm by Chang et al. (ChangJP19, ) described above.
@@
Subsequently, this line of work has been extended to several challenging settings, showing constant competitiveness so long as $\gamma$ is known to all the participants, and the algorithm is executed for a sufficiently large number of slots. Richa et al. (richa:jamming2, ) consider a multi-hop network modeled as a unit disk graph (UDG). A description of the UDG has been given by Schmid and Wattenhofer (schmid2006algorithmic, ). First, Richa et al. (richa:jamming2, ; DBLP:journals/dc/RichaSSZ13, ) consider the setting where the adversary is 1-uniform and the UDG is connected, or each participant has at least $2/\epsilon$ neighbors. Then, against a $1$-uniform $(T,1-\epsilon)$-bounded adversary, the authors provide an algorithm, named Jade, that is w.h.p. constant competitive. Second, Richa et al. (richa:jamming2, ) prove that, if the graph is not connected, or the adversary is $2$-uniform—for any slot, the adversary may choose any partition of the $n$ participants and have the members of one partition hear jamming, while the others do not—and participants each have $o(1/\epsilon)$ neighbors, then Jade can fail to be constant competitive.
Against a reactive jamming adversary—who can learn whether or not the channel is idle in the current slot and can choose to jam based on this information—Richa et al. (richa:jamming3, ; DBLP:journals/ton/RichaSSZ13, ) show that constant competitiveness is still possible. In addition to jamming, more benign noise, such as interference due to environmental factors, can present a challenge to communication.
@@
Under the SINR model (recall Section 3.3.2), Ogierman et al. (ogierman:competitive, ; ogierman2018sade, ) address jamming. The authors provide a multiplicative weights update algorithm that is $2^{-O(1/\epsilon)^{2/\alpha-2}}$-competitive, where $\alpha>2$ is a parameter of the SINR model known as the path-loss exponent, which corresponds to the degree of signal attenuation.
@@
A different approach is taken by Bender et al. (bender:how, ; BenderFGY19, ), who address an adversary who is not constrained in its jamming. A packet starts in an “inactive” state by listening to the channel, and if a sufficient number of empty slots are heard, the packet becomes “active” in the sense of making transmission attempts. This “throttling” mechanism gives time for contention to decrease, thus allowing active packets to succeed. An active packet will make transmission attempts, but in contrast to updating based on feedback in each slot, Bender et al. (bender:how, ; BenderFGY19, ) monotonically reduce the sending probability of each packet $u$ as roughly $1/s_{u}$, where $s_{u}$ is the number of slots since $u$ most recently started attempting to transmit (i.e., the age of $u$). Packet also listens (when not transmitting), and if a sufficient number of empty slots are heard, it resets $s_{u}=1$. Here, they show expected constant throughput, with each packet making an expected number of transmission attempts that is polylogarithmic in $n+J$, where $J$ is the total number of jammed slots during which there exists at least one packet. An analogous result is also given for the case of infinite packets, illustrating constant throughput and an expected average number of transmission attempts that is polylogarithmic in $\eta_{t}+J_{t}$, where $\eta_{t}$ is the maximum number of concurrently sent packets in the system before slot $t$, and $J_{t}$ is the number of jammed slots before slot $t$.
For the wakeup problem (recall Section 3.3.2), Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) consider a multi-channel setting where there are $k$ out of $n$ participants activated, each of which may access $b$ channels simultaneously in any round. The adversary is a random jammer: each channel is jammed independently and uniformly with known probability $p<1$ per round. The authors present two deterministic algorithms for unknown $k$ that succeed with probability $1-O(1/\texttt{poly}(n))$, the first with latency $O(k\log(n)\log^{1/b}k/\log(1/p))$, and the second with latency $O((k/b)\log(n)\log(b\log(n))/\log(1/p))$ when $b=\Omega(\log(b\log n))$.
@@
The AQT setting (recall Section 3.1) has been extended to incorporate jamming. Anantharamu et al. (anantharamu2019packet, ; anantharamu2011medium, ) extends a leaky-bucket adversary (recall Section 3.1) to ${(\lambda,\rho,T)}$-bounded leaky bucket adversary who can also jam $\lambda t+b$ slots in any window of consecutive $t$ slots; here, $\lambda$ is the jamming rate, while $b$ is the burstiness parameter (recall Section 3.1). Deterministic algorithms are presented, with and without CD, giving a range of bounds on queue size and latency parameterized by the injection rate, burstiness, and the known number of stations $n$, each of which has a unique ID in the range $[0,...,n-1]$.
Similarly, Anta et al. (DBLP:journals/tcs/AntaGKZ17, ) consider an adversary whose jamming capability is parameterized by both a rate and the number of packets it can jam. This adversary is referred to as ($\sigma,\rho$)-bounded, where $\rho\in[0,1]$ and $\sigma\geq 1$. The adversary has a token bucket of size $\sigma$ where it stores error tokens. Adversary can introduce an error in the channel and jam the packet being transmitted, consuming one token. For any time $t$, within interval $T=[0,t]$, there will be $\lfloor\rho T\rfloor$ created, where $\rho$ is the rate at which they become available to the adversary. Here, the authors consider deterministic algorithms for transmitting data from a single sender to a single receiver. The adversary is assumed to know the algorithm and use this information to inform its jamming strategy. The amount of data (i.e., packet size) may be adjusted by the sender during execution. If the channel is jammed during transmission, the sender immediately becomes aware of this failure.
@@
In constrast to the results discussed above, Jurdzinski et al. (DBLP:conf/waoa/JurdzinskiKL14, ) consider an unbounded adversary. Packets of different sizes may arrive over time. If the adversary jams at any time during which a packet is transmitted, the transmission fails and the sender learns of this immediately. Under this model, the authors give a deterministic, constant-competitive online algorithm.
@@
An impossibility result is given by Bender et al. (bender2020contention, ), who show that, without CD, no algorithm can guarantee a packet success in time linear in the system size with better than constant probability in the presence of jamming; this result holds even if against a non-adaptive jamming adversary. Subsequently, Chen et al. (DBLP:conf/podc/ChenJZ21, ) give an algorithm that achieves a trade-off between the throughput and the amount of jamming that can be tolerated. In the challenging case where a constant fraction of slots are jammed, the authors achieve $O(1/\log n)$ throughput, which aligns with the lower bound of Bender et al. (bender2020contention, ).
Discussion.
@@
Perhaps unsurprisingly, randomization along with CD can provide constant throughput (in non-jammed slots) despite an unbounded adversary; even without CD, jamming can be tolerated, although throughput decreases slightly to $\Theta(1/\log n)$. When the adversary is bounded, a range of results exist. Under a saturated model, randomized results demonstrate constant throughput for a variety of network settings against a bounded adversary, so long as some weak information about the adversary’s power and $n$ is known. In contrast, deterministic results in the AQT setting are more nuanced, being parameterized by aspects of jamming rate and burstiness, and generally more pessimistic; however, strong online results are possible.
@@
Existing results that tolerate jamming exhibit a sharp trade-off between the amount of listening devices perform and throughput achieved. On one hand, if each participant listens to the channel in every slot (except when sending), then constant throughput (in expectation or w.h.p. in non-jammed slots) is possible (bender:how, ; BenderFGY19, ; ChangJP19, ; awerbuch:jamming, ; richa:jamming2, ; richa:jamming3, ). Note that such results are not energy efficient, given that listening is roughly as costly as sending in practice (recall Section 1.1.3). On the other hand, if devices never listen to the channel, only sub-constant throughput is possible, and the best result is $O(1/\log n)$ throughput w.h.p. (DBLP:conf/podc/ChenJZ21, ). Therefore, an open problem for jamming-resistant contention resolution is whether constant throughput can be achieved with a small amount of listening.

### v3

@@ 4. Jamming-Resistant Contention Resolution
There are a number of papers addressing applied security considerations with respect to jamming (for examples, see (bayraktaroglu:performance, ; xu:feasibility, ; deejam:wood, ; aschenbruck:simulative, ; wood:denial, ; xu:jamming, ) and surveys (mpitziopoulos:survey, ; VADLAMANI201676, )). For example, spread spectrum techniques (spread:foiling, ; navda:using, ), rerouting defenses (wood2003jam, ; han2018hierarchical, ), and specialized antenna technology (zheng2018design, ; mcmilin2015gps, ). From a theory perspective, several fundamental distributed computing challenges—such as broadcast (koo2, ; gilbert:near, ; gilbert:making, ; chen:broadcasting, ), leader election (gilbert:malicious, ; gilbert2009malicious, ), gossip (gilbert:interference, ; dolev:gossiping, ), node discovery (meier:speed, ), and simple point-to-point communication (king:conflict, ; DBLP:journals/dc/KingPSY18, )—can be solved on a shared channel despite jamming (see (young:overcoming, ; Bender:2015:RA:2818936.2818949, )). Here, we focus on prior contention-resolution results that offer provable guarantees on throughput or latency, despite jamming.
@@
A jamming adversary poses at least two significant new challenges over the dynamic-arrivals model considered in Section 3. First, it may not be obvious that non-trivial makespan guarantees are possible; for example, if jamming occurs in every slot in perpetuity, then no participants can succeed. Therefore, we must either constrain the jamming power of the adversary (e.g., by assuming a bounded adversary), or alter our measure of throughput to address this case (e.g., such as notions of competitiveness and implicit throughput). Second, the algorithms surveyed so far that rely on channel feedback to adjust contention—for example, by estimating $n$ (such as in (BenderKPY18, ; DBLP:journals/jacm/GreenbergFL87, )) or by exchanging messages between participants—are vulnerable, since jamming allows the adversary to create collisions and prevent communication.
With regard to the first challenge, a common approach is to measure performance over non-jammed slots; for example, this is done by Chang et al. (ChangJP19, ) and Bender et al. (bender:how, ). Specifically, assuming that at least one participant is present in the system, a slot that is not jammed, but where no participant succeeds, is said to be wasted. Guaranteeing that only a small constant fraction of slots are wasted allows us to derive meaningful results despite an adversary that jams aggressively, perhaps even in every slot.
Addressing the second challenge of adjusting the contention requires care, and this is demonstrated by the elegant algorithm by Chang et al. (ChangJP19, ). Participants follow some simple rules for updating the sending probability in each slot: increasing if there is silence, decreasing if there is noise, and leaving it unchanged in the case of a success. The adversary may lower the sending probabilities of participants by jamming (i.e., adding noise), thus depressing contention. However, the update rule allows for contention to quickly recover once the jamming ceases (although, it may resume again later), and few slots are wasted as a result. Notably, this design allows the algorithm to achieve roughly $1/e$ throughput in the non-jammed slots, which rivals that of slotted ALOHA, despite the more-challenging adversarial setting.
@@
An early series of results against a bounded adversary was initiated by Awerbuch et al. (awerbuch:jamming, ; DBLP:journals/talg/AwerbuchRSSZ14, ), who consider the following model. There are $n$ honest participants equipped with CD, and the network is saturated (recall Section 2.1). For some $T\in\mathbb{N}$ and constant $10<\epsilon<$ 1, there is an adversary that can jam at most $(1-\epsilon)w$ of the slots in any consecutive $w\geq T$ slots. This adversary is referred to as $(T,1-\epsilon)$-bounded, and the value of $\epsilon$ is unknown to participants.
@@
Such bounded jamming may be well-motivated in practice, where aggressive jamming can reveal that an attack is occurring, as well as the location of the malicious device(s) (for examples, see (liu2009localizing, ; liu2013error, ; xu:feasibility, )), or correspond to a limited power supply. The adversary is also adaptive, meaning that the adversary knows the full history of the actions performed by honest nodes for all the previous slots, and may use this information to inform its jamming strategy in the current slot; however, the adversary does not know what actions participants are taking in the current slot. When the adversary jams a slot, all listening participants hear jamming; this is referred to as a 1-uniform adversary.
@@
Awerbuch et al. (awerbuch:jamming, ; DBLP:journals/talg/AwerbuchRSSZ14, ) define an algorithm to be constant competitive if, for any sufficiently large number of slots, the honest nodes perform successful message transmissions in a constant fraction of the non-jammed steps. The authors provide a randomized algorithm that is constant competitive with high probability (w.h.p.) in $n$, when all participants know a common parameter, $\gamma$$=O(1/(\log T+\log\log n))$, and the algorithm is run for a sufficiently large number of slots. Roughly speaking, in a slot, the algorithm dictates that each participant $u$ sends a message with probability $p_{u}$. If $u$ does not send, it listens and updates $p_{u}$ via a multiplicative factor which can increase or decrease, depending on whether or not the slot is idle. Notably, this multiplicative weights update method shares common ground with the later algorithm by Chang et al. (ChangJP19, ) discussed above.
@@
Subsequently, this line of work has been extended to several challenging settings, showing constant competitiveness so long as $\gamma$ is known to all the participants, and the algorithm is executed for a sufficiently large number of slots. Richa et al. (richa:jamming2, ) consider a multi-hop network modeled as a unit disk graph (UDG). Informally, under the UDG model, each participant can broadcast within a disk centered about its location, and two participants may communicate if their respective disks intersect (see Schmid and Wattenhofer (schmid2006algorithmic, ) for discussion of the UDG model). First, Richa et al. (richa:jamming2, ; DBLP:journals/dc/RichaSSZ13, ) consider the setting where the adversary is 1-uniform and the UDG is connected, or each participant has at least $2/\epsilon$ neighbors. Then, against a $1$-uniform $(T,1-\epsilon)$-bounded adversary, the authors provide an algorithm, named Jade, that is w.h.p. constant competitive. Second, Richa et al. (richa:jamming2, ) prove that, if the graph is not connected, or the adversary is $2$-uniform—for any slot, the adversary may choose any partition of the $n$ participants and have the members of one partition hear jamming, while the others do not—and participants each have $o(1/\epsilon)$ neighbors, then Jade can fail to be constant competitive. Against a reactive jamming adversary—who can learn whether or not the channel is idle in the current slot and can choose to jam based on this information—Richa et al. (richa:jamming3, ; DBLP:journals/ton/RichaSSZ13, ) show that constant competitiveness is still possible.
@@
Under the SINR channel model (recall Section 2.3), Ogierman et al. (ogierman:competitive, ; ogierman2018sade, ) address jamming in a more-general network topology than prior results on competitiveness. The authors provide a multiplicative weights update algorithm that is $2^{-O(1/\epsilon)^{2/(\alpha-2)}}$-competitive, where $\alpha>2$ is a parameter of the SINR model known as the path-loss exponent, which corresponds to the degree of signal attenuation. This competitiveness holds if the algorithm is executed for a sufficiently large number of slots against a $1$-uniform adversary and the transmission range contains at least one node, or at least $2/\epsilon$ nodes within the transmission range of every node.
@@
A different approach is taken by Bender et al. (bender:how, ; BenderFGY19, ), who address an adversary that is not constrained in its jamming. A packet starts in an “inactive” state by listening to the channel, and if a sufficient number of empty slots are heard, the packet becomes “active” in the sense of making transmission attempts. This “throttling” mechanism gives time for contention to decrease, thus allowing active packets to succeed. An active packet $u$ will make transmission attempts, but in contrast to updating based on feedback in each slot, $u$ monotonically reduces its sending probability as roughly $1/s_{u}$, where $s_{u}$ is the number of slots since $u$ became active (i.e., the age of $u$). Packet $u$ also listens (when not transmitting), and if a sufficient number of empty slots are heard, it resets $s_{u}=1$. Here, they show expected constant throughput, with each packet making an expected number of transmission attempts that is polylogarithmic in $n+J$, where $J$ is the total number of jammed slots during which there exists at least one packet.
Recently, Bender et al. (bender:fully, ) showed how to achieve $\Omega(1)$ implicit throughput (recall Section 1.1.3) with $O({\texttt{poly}({\log n})})$ channel accesses (both sending and listening), despite an adaptive jamming adversary. This result applies in the channel setting with CD (i.e., ternary channel feedback), but with no control bits allowed, unlike the energy-efficient result in (BenderKPY18, ). The algorithm, named Low-Sensing Backoff, is based on multiplicative weights updates, and in that regard, it shares common ground with some prior results (such as (ChangJP19, ; richa:jamming2, ; richa:jamming3, ; richa:jamming4, ; awerbuch:jamming, ; ogierman2018sade, )). However, in comparison to these results, Low-Sensing Backoff greatly limits the frequency at which packets can sense the channel in order to update, while the lack of control bits greatly limits the amount of information they receive in each update. Analogous results are given for the cases where the number of packets is infinite, and where the adversary is reactive.
@@
The AQT setting (recall Section 3.1) has been extended to incorporate jamming. Anantharamu et al. (anantharamu2019packet, ; anantharamu2011medium, ) extends a leaky-bucket adversary to ${(\lambda,\rho,T)}$-bounded leaky bucket adversary who can also jam $\lambda t+b$ slots in any window of consecutive $t$ slots; here, $\lambda$ is the jamming rate, while $b$ is the burstiness parameter. Deterministic algorithms are presented, with and without CD, giving a range of bounds on queue size and latency parameterized by the injection rate, burstiness, and the known number of stations $n$, each of which has a unique ID in the range $[0,...,n-1]$.
In a related setting, but with a single sender and receiver, Anta et al. (DBLP:journals/tcs/AntaGKZ17, ) consider an adversary whose jamming capability is parameterized by a rate, $\rho\in[0,1]$, and the length of the largest burst of jammed packets, $\sigma\geq 1$; both parameters are known to the algorithm. Each packet has a fixed-length header, along with a payload whose size may be adjusted by the sender. The authors derive a lower bound on the optimal rate of transmission for payload information (i.e., goodput), and also develop an adaptive deterministic algorithm that is optimal.
@@
Jurdziński et al. (DBLP:conf/waoa/JurdzinskiKL14, ) also consider a single sender and receiver, but in contrast to a parameterized adversary, they consider: an unbounded adversary, and packets of different sizes may arrive over time. If the adversary jams at any time during which a packet is transmitted, the transmission fails and the sender learns of this immediately. The authors derive results on the ratio between the throughput achieved by the algorithm and the throughput achieved by the best scheduling under identical arrival and error patterns; this is referred to as relative throughput. The authors give a deterministic online algorithm that achieves optimal relative throughput.
@@
An impossibility result is given by Bender et al. (bender2020contention, ), who show that with no-CD, no algorithm can guarantee a packet success in time linear in the system size with better than constant probability in the presence of jamming; this result holds even if against a non-adaptive adversary. Subsequently, Chen et al. (DBLP:conf/podc/ChenJZ21, ) give an algorithm in the no-CD case that achieves a trade-off between the throughput and the amount of jamming tolerated. In the challenging case where a constant fraction of slots are jammed, $O(1/\log n)$ throughput is shown, which aligns with the lower bound of Bender et al. (bender2020contention, ).
@@
Summary. Perhaps unsurprisingly, randomization along with CD can provide constant throughput (in non-jammed slots) despite an unbounded adversary; even without CD, jamming can be tolerated, although throughput decreases slightly to $\Theta(1/\log n)$. When the adversary is bounded, a range of results exist. Under a saturated model, randomized results demonstrate constant throughput for a variety of network settings against a bounded adversary, so long as some weak information about the adversary’s power and $n$ is known. In contrast, deterministic results in the AQT setting are more nuanced, being parameterized by aspects of jamming rate and burstiness, and generally more pessimistic.
@@
Existing results that tolerate jamming exhibit a sharp trade-off between the amount of listening devices perform and throughput achieved. On one hand, if each participant listens to the channel in every slot (except when sending), then constant throughput (in expectation or w.h.p. in non-jammed slots) is possible (bender:how, ; BenderFGY19, ; ChangJP19, ; awerbuch:jamming, ; richa:jamming2, ; richa:jamming3, ). Note that such results are not energy efficient, given that listening is roughly as costly as sending in practice (recall Section 1.1.3). On the other hand, if devices never listen to the channel, only sub-constant throughput is possible, and the best result is $O(1/\log n)$ throughput w.h.p. (DBLP:conf/podc/ChenJZ21, ). The recent result in (bender:fully, ) demonstrates that jamming-resistant contention resolution is possible with constant throughput and a polylogarithmic number of channel accesses (both sending and listening) without using control bits. However, it is unclear whether better energy-efficiency bounds are possible in the presence of jamming, perhaps under models with additional feedback permitted on the channel.